# On the analysis of a higher-order Lotka-Volterra model: an application of $\mathcal{S}$-tensors and the polynomial complementarity problem

### v1

@@ VI Purely competitive higher-order Lotka-Volterra
where $a_{ij}$ and $b_{ijk}$ are all non-negative. In this scenario, all species have no friends and they all compete with each other.
Lemma 12
@@
Proof:
It is sufficient to check that $\dot{x}_{i}=0$ when $x_{i}=0$ and $\dot{x}_{i}=r_{i}x_{i}\left(1-\sum^{n}_{j=1}a_{ij}x_{j}-\sum^{n}_{j,k=1}b_{ijk}x_{j}x_{k}\right)\leq r_{i}x_{i}-r_{i}a_{ii}x_{i}^{2}=r_{i}x_{i}(1-a_{ii}x_{i})<0$ when $x_{i}>\frac{1}{a_{ii}}$. ∎
@@
The existence of at least one equilibrium is guaranteed by Corollary 2. We recall that a strictly diagonally dominant tensor with positive diagonal elements is a $\mathcal{S}$-tensor, see [15]. For example, we consider $A$ and $B$ are strictly diagonally dominant tensors, since $-A$ and $-B$ are with positive diagonal elements, then $-A$ and $-B$ are $\mathcal{S}$-tensors. The Corollary 2 is thus directly applicable. Once there is a positive equilibrium, the positive equilibrium is globally attractive when the condition of corollary 6 is fulfilled. Otherwise, if $A,B$ are both strictly diagonally dominant tensors with negative diagonal entries, then it is globally asymptotically stable according to the Theorem 5. It is worthwhile to mention that in the case of pure competition, we don’t need the assumption of existence of a upper bound in corollary 6. The general upper bound in the case is $R=max_{i}\frac{1}{a_{ii}}$. This leads to the following result:
Corollary 10
@@
Consider the system (37). Let $\hat{R}=max_{i}\frac{1}{a_{ii}}$. If there holds $-d_{i}a_{ii}>\sum_{j\neq i}d_{j}|a_{ij}|+\hat{R}\sum_{j\neq i,1\leq k\leq n}(d_{j}|b_{ijk}|+d_{j}|b_{ikj}|)$ for all $i$, then the positive equilibrium $x^{*}$, if it exists, is unique and globally asymptotically stable.
@@
While the positive equilibrium corresponds to the winner-share-all consequence, the following result is about the winner-take-all case. This result is an extension of Corollary 7.4 [36] and Theorem 15 [4] for a classical competitive Lotka-Volterra model.
Theorem 16
@@
Consider the system (37) and the following conditions: (A) $\frac{1}{a_{jj}}<\frac{1}{a_{ij}},\quad 1\leq i<j\leq n$ and $(B)\frac{1}{a_{jj}}>\frac{1}{a_{ij}},n\geq i>j\geq 1$. Then, $\left(\frac{-a_{11}+\sqrt{a_{11}^{2}+4b_{111}}}{2b_{111}},0,\ldots,0\right)$ is globally attracting on $\mathbb{R}_{n}^{+}$.
Proof:
@@
The higher-order system (37) is upper-bounded by a classical competitive Lotka-Volterra counterpart (chapter 5 [4]). We know that $\left(\frac{1}{a_{11}},0,\ldots,0\right)$ is globally attracting on $\mathbb{R}_{n}^{+}$ for a classical competitive Lotka-Volterra under the conditions (A) and (B) (Theorem 15 [4]). If $i\neq 1$, then $x_{i}$ will converge to zero since the classical model converges to zero. If $i=1$, then $\dot{x_{i}}$ is an autonomous system $\dot{x_{i}}=1-a_{ii}x_{i}-b_{iii}x_{i}^{2}$ plus a vanishing perturbation. The solution of $\dot{x_{i}}=1-a_{ii}x_{i}-b_{iii}x_{i}^{2}$ will converge to $\left(\frac{-a_{11}+\sqrt{a_{11}^{2}+4b_{111}}}{2b_{111}},0,\ldots,0\right)$. We then see that $\frac{\frac{-a_{11}+\sqrt{a_{11}^{2}+4b_{111}}}{2b_{111}}}{\frac{1}{a_{11}}}=\frac{a_{11}}{a_{11}+\sqrt{a_{11}^{2}+4b_{111}}}\leq 1$. This stable equilibrium is also upper-bounded by the equilibrium of its classical competitive Lotka-Volterra counterpart.
Thus, the solution of the system (37) will converge to $\left(\frac{-a_{11}+\sqrt{a_{11}^{2}+4b_{111}}}{2b_{111}},0,\ldots,0\right)$. ∎
Remark 15
Notice that the Theorem 16 describes the case when the first species is the winner. However, via a simple index permutation, we can further know the case when the other species is the winner.
@@
Theorem 17
Consider the system (37), the boundary equilibrium $M^{*}=(x_{1}^{*},\cdots,x_{a}^{*},\mathbf{0}_{n-a})$, if it exists, is locally stable when the coefficients of the first- ($a_{ij},i\neq j$) and second-order competitive terms ($d_{ijk},j,k\in Q_{i}$) of the losers are sufficiently large such that $01-\sum_{j\in S}a_{ij}x^{*}_{j}-\sum_{j,k\in S}c_{ijk}x^{*}_{j}x^{*}_{k}<$ 0, and moreover $(x_{1}^{*},\cdots,x_{a}^{*})$ is a stable positive equilibrium point of the sub-system of the winners when ignoring the loser agents.
Proof:
The proof is similar to the proof of Theorem 15. ∎
Remark 16
Since there are multiple sub-systems, each subsystem may have a stable positive equilibrium. In general, multi-stability may occur.
VII Numerical example
@@
Firstly, we conduct simulations for the case of competition between 2 factions. We assume a total of $5$ species, $2$ in one faction and $3$ in the other. Figures 1(a)-1(c) include all the possible results we can observe and are three typical examples for each case. Figure 1(a) is in line with Theorem 12, while Figure 1(b) corresponds to Theorem 13. Similarly, Figure 1(c) reflects the analytical results of Theorem 15. Since all other equilibria are unstable, we don’t observe that the solution converges to them from in simulation scenarios. From figure 1(a) and 1(d), we confirm the bistability properties of the two-faction system. Furthermore, to highlight the influence of HOIs, the system parameters and initial conditions are the same in simulations 1(b) and 1(c) except that we change $d_{1jk}$ to $5d_{1jk}$, respectively. As we increase the $d_{1jk}$, according to the Theorem 15, the first faction dies out, which is indeed in line with the simulation results. This shows that our theory can also be seen as a manipulation strategy to adjust the winner species. Since HOIs usually denote the indirect interaction in ecology, HOIs are potentially more suitable to adjust than pairwise direct interaction. Finally, from simulations, we observe that large self-competition terms ($a_{ii},\hat{a}_{ii},c_{iii},\hat{c}_{iii}$) will improve the chance that the system doesn’t diverge to infinity. This is because diagonally dominant tensors with all positive diagonal entries are $\mathcal{S}$-tensors, which potentially yields a globally stable positive equilibrium.
@@
Then, we use a more general case. For the tensor $A$, its entries are $A_{111}=A_{222}=11,A_{112}=A_{122}=A_{211}=A_{212}=-1,A_{121}=10,A_{221}=5$. For the matrix $B$, $B_{11}=0.8883,B_{12}=1,B_{21}=0,B_{22}=1$. Let $v=(0.1117,1)^{\top}$. We can check that $Av^{2}>\mathbf{0}$ and $Bv>\mathbf{0}$. The tensors $A,B$ are all $\mathcal{S}-$tensors but not $\mathcal{H}^{+}-$ tensors. Thus, the conditions of Theorem 1 are satisfied. Further simulations suggest that $Ax^{2}+Bx=\mathbf{1}$ has a unique positive solution $(0.49547,0.43791)^{\top}.$
VIII Conclusion
This paper proposes a general higher-order Lotka-Volterra model. More precisely, we consider 3 typical scenarios: full cooperation, competition between 2 factions, and pure competition respectively. For the analysis tool, we provide an existence result of a positive equilibrium for a non-homogeneous polynomial system and give a lower and upper bound for the solution of a polynomial complementarity problem under mild conditions, which is an extension of the current results in tensor algebras. Then, we further utilize the properties of tensor and monotone system theory to provide the results regarding the existence, uniqueness, and stability of the equilibrium of the system. This paper yields a comprehensive understanding of a higher-order Lotka-Volterra model. Finally, all theoretical results are illustrated by some numerical examples.

### v2

@@ 6 Purely competitive higher-order Lotka-Volterra
with $i=1,\ldots,n$, and where $a_{ij}$ and $b_{ijk}$ are all non-negative. In this scenario, all species have no friends and they all compete with each other.
Lemma 6.1.
@@
Proof 6.2.
It is sufficient to check that $\dot{x}_{i}=0$ when $x_{i}=0$ and $\dot{x}_{i}=r_{i}x_{i}\left(1-\sum^{n}_{j=1}a_{ij}x_{j}-\sum^{n}_{j,k=1}b_{ijk}x_{j}x_{k}\right)\leq r_{i}x_{i}-r_{i}a_{ii}x_{i}^{2}=r_{i}x_{i}(1-a_{ii}x_{i})<0$ when $x_{i}>\frac{1}{a_{ii}}$.
@@
The existence of one unique positive equilibrium is guaranteed by Theorem 11. We recall that a strictly diagonally dominant tensor with positive diagonal elements is a $\mathcal{S}$-tensor, see [15]. For example, we consider $A$ and $B$ are strictly diagonally non-positive dominant tensors and $-A$ and $-B$ are with positive diagonal elements. Then $-A$ and $-B$ are $\mathcal{S}$-tensors. The Theorem 11 is thus directly applicable. Once there is a positive equilibrium, the positive equilibrium is globally attractive when the condition of corollary 9 is fulfilled. It is worthwhile to mention that in the case of pure competition, we don’t need the assumption of the existence of an upper bound in corollary 9. The general upper bound in the case is $R=max_{i}\frac{1}{a_{ii}}$. This leads to the following result:
Corollary 23.
@@
Consider the system (34). Let $\hat{R}=\max_{i}\frac{1}{a_{ii}}$. If there holds $-d_{i}a_{ii}>\sum_{j\neq i}d_{j}|a_{ij}|+\hat{R}\sum_{j\neq i,1\leq k\leq n}(d_{j}|b_{ijk}|+d_{j}|b_{ikj}|)$ for all $i$, then the positive equilibrium $x^{*}$, if it exists, is unique and globally asymptotically stable.
Otherwise, the existence and stability can be checked with the following result, which is a stronger result.
Theorem 24.
System (20) under the competitive setting (i.e., (34)) has one unique positive equilibrium if $-B$ is an irreducible nonnegative $\mathcal{H}^{+}$-tensor, $-A$ is an irreducible nonnegative $\mathcal{H}^{+}$-matrix. The equilibrium point $x^{*}$ is globally asymptotically stable.
Proof 6.3.
Notice that an irreducible nonnegative tensor is a $\mathcal{S}$-tensor and for any $v>\mathbf{0}$, $Av>\mathbf{0}$ and $Bv^{2}>\mathbf{0}$. Thus, the system has a unique positive equilibrium according to Theorem 11. Define the set $U_{1}=\{x|x\geq x^{*}\}$ and $U_{2}=\{x|\min\frac{x_{i}}{x_{i}^{*}}<1\}$.
Then, we let $V_{m}=\min_{i}(\frac{x_{i}}{x^{*}_{i}})^{2}$ and $V=\min_{i}(\frac{x_{i}-x_{i}^{*}}{x^{*}_{i}})^{2}$. Further let $m=\arg\min_{i}(\frac{x_{i}}{x^{*}_{i}})^{2}$ and $t=\arg\min_{i}(\frac{x_{i}-x^{*}_{i}}{x^{*}_{i}})^{2}$. It holds $m=t$ because of $\frac{x_{i}-x^{*}_{i}}{x^{*}_{i}}=\frac{x_{i}}{x^{*}_{i}}-1$.
Furthermore, it must hold $-x_{i}=-\frac{x_{i}x_{i}^{*}}{x_{i}^{*}}\leq-V_{m}^{\frac{1}{2}}x_{i}^{*}.$ It holds as a equality when $i=m$, otherwise it holds as an strict inequality.
Then, we have in the set $U_{1}$
$\begin{split}\dot{V}&=\frac{2r_{m}x_{m}}{(x_{m}^{*})^{2}}(x_{m}-x_{m}^{*})(Ax+Bx^{2}+\mathbf{1})_{m}\\ &=\frac{2r_{m}x_{m}}{(x_{m}^{*})^{2}}(x_{m}-x_{m}^{*})(\sum_{j}a_{mj}x_{j}+\sum_{jk}b_{mjk}x_{j}x_{k}+1)\\ &\leq\frac{2r_{m}x_{m}}{(x_{m}^{*})^{2}}(x_{m}-x_{m}^{*})(\sum_{j}a_{mj}x_{j}^{*}V_{m}^{\frac{1}{2}}+\sum_{jk}b_{mjk}x^{*}_{j}x^{*}_{k}V_{m}+1)\\ &\leq\frac{2r_{m}x_{m}}{(x_{m}^{*})^{2}}(x_{m}-x_{m}^{*})(\sum_{j}a_{mj}x_{j}^{*}V_{m}+\sum_{jk}b_{mjk}x^{*}_{j}x^{*}_{k}V_{m}+1)\\ &=\frac{2r_{m}x_{m}}{(x_{m}^{*})^{2}}(x_{m}-x_{m}^{*})(-V_{m}+1)\leq 0\quad\forall x\in U_{1}.\end{split}$
Otherwise, for the set $U_{2}$
$\begin{split}\dot{V}&\leq\frac{2r_{m}x_{m}}{(x_{m}^{*})^{2}}(x_{m}-x_{m}^{*})(\sum_{j}a_{mj}x_{j}^{*}V_{m}^{\frac{1}{2}}+\sum_{jk}b_{mjk}x^{*}_{j}x^{*}_{k}V^{\frac{1}{2}}_{m}+1)\\ &=\frac{2r_{m}x_{m}}{(x_{m}^{*})^{2}}(x_{m}-x_{m}^{*})(-V^{\frac{1}{2}}_{m}+1)\leq 0\quad\forall x\in U_{2}.\end{split}$
Notice that $V$ is locally positive definite in $U_{1}$ and $U_{2}$, and $\dot{V}$ is negative semidefinite. The solution must converge to $\{x|\dot{V}=0\}$. In order to have $\dot{V}=0$, for every $i$, it must hold $i=m$ and $x_{m}=x_{m}^{*}$ which indicates that $\{x|\dot{V}=0\}$ only contains $x^{*}$. Thus, $x^{*}$ is asymptotically stable with a domain of attraction $U_{1},U_{2}$.
Next, we further let $L_{s}=\max_{i}(\frac{x_{i}}{x^{*}_{i}})^{2}$ and $s=\arg\max_{i}(\frac{x_{i}}{x^{*}_{i}})^{2}$.
Furthermore, it must hold $-x_{i}=-\frac{x_{i}x_{i}^{*}}{x_{i}^{*}}\geq-L_{s}^{\frac{1}{2}}x_{i}^{*}.$ It holds as a equality when $i=s$, otherwise it holds as an strict inequality.
Then, consider the set $U_{3}=\{x|x<x^{*}\}$, we have
$\begin{split}\dot{V}&=\frac{2r_{s}x_{s}}{(x_{s}^{*})^{2}}(x_{s})(Ax+Bx^{2}+\mathbf{1})_{s}\\ &=\frac{2r_{s}x_{s}^{2}}{(x_{s}^{*})^{2}}(\sum_{j}a_{sj}x_{j}+\sum_{jk}b_{sjk}x_{j}x_{k}+1)\\ &\geq\frac{2r_{s}x_{s}^{2}}{(x_{s}^{*})^{2}}(\sum_{j}a_{sj}x_{j}^{*}L_{m}^{\frac{1}{2}}+\sum_{jk}b_{sjk}x^{*}_{j}x^{*}_{k}L_{m}+1)\\ &\geq\frac{2r_{s}x_{s}^{2}}{(x_{s}^{*})^{2}}(\sum_{j}a_{sj}x_{j}^{*}L_{m}+\sum_{jk}b_{sjk}x^{*}_{j}x^{*}_{k}L_{m}+1)\\ &=\frac{2r_{s}x_{s}^{2}}{(x_{s}^{*})^{2}}(-L_{m}+1)>0\quad\forall x\in U_{3}.\end{split}$
This guarantees that all trajectories in $U_{3}$ finally enter $U_{1}\cup U_{2}$. Thus, the $x^{*}$ is globally asymptotically stable.
@@
While the positive equilibrium corresponds to the winner-share-all consequence, the following result is about the winner-take-all case. This result is an extension of [38, Corollary 7.4] and [4, Theorem 15] for the classical competitive Lotka-Volterra model.
Theorem 25.
Consider the system (34) and the following conditions:
  • (A)
    $\frac{1}{a_{jj}}<\frac{1}{a_{ij}},\quad 1\leq i<j\leq n$, and
  • (B)
    $\frac{1}{a_{jj}}>\frac{1}{a_{ij}},n\geq i>j\geq 1$.
@@
Then, $\left(\frac{-a_{11}+\sqrt{a_{11}^{2}+4b_{111}}}{2b_{111}},0,\ldots,0\right)$ is globally attracting on $\mathbb{R}_{n}^{+}$.
Proof 6.4.
@@
The higher-order system (34) is upper-bounded by a classical competitive Lotka-Volterra counterpart ([4, chapter 5]). We know that $\left(\frac{1}{a_{11}},0,\ldots,0\right)$ is globally attracting on $\mathbb{R}_{n}^{+}$ for a classical competitive Lotka-Volterra under the conditions (A) and (B) ([4, Theorem 15]). If $i\neq 1$, then $x_{i}$ will converge to zero since the classical model converges to zero. If $i=1$, then $\dot{x_{i}}$ is an autonomous system $\dot{x_{i}}=1-a_{ii}x_{i}-b_{iii}x_{i}^{2}$ plus a vanishing perturbation. The solution of $\dot{x_{i}}=1-a_{ii}x_{i}-b_{iii}x_{i}^{2}$ will converge to $\left(\frac{-a_{11}+\sqrt{a_{11}^{2}+4b_{111}}}{2b_{111}},0,\ldots,0\right)$. We then see that $\frac{\frac{-a_{11}+\sqrt{a_{11}^{2}+4b_{111}}}{2b_{111}}}{\frac{1}{a_{11}}}=\frac{a_{11}}{a_{11}+\sqrt{a_{11}^{2}+4b_{111}}}\leq 1$. This stable equilibrium is also upper-bounded by the equilibrium of its classical competitive Lotka-Volterra counterpart.
Thus, the solution of the system (34) will converge to $\left(\frac{-a_{11}+\sqrt{a_{11}^{2}+4b_{111}}}{2b_{111}},0,\ldots,0\right)$.
Remark 6.5.
Notice that Theorem 25 describes the case when the first species is the winner. However, via a simple index permutation, we can also describe the case when the other species is the winner.
@@
Theorem 26.
Consider the system (34). The boundary equilibrium $M^{*}=(x_{1}^{*},\cdots,x_{a}^{*},\mathbf{0}_{n-a})$, if it exists, is locally stable when the coefficients of the first- ($a_{ij},i\neq j$) and second-order competitive terms ($d_{ijk},j,k\in Q_{i}$) of the losers are sufficiently large such that $01-\sum_{j\in S}a_{ij}x^{*}_{j}-\sum_{j,k\in S}c_{ijk}x^{*}_{j}x^{*}_{k}<$ 0, and moreover $(x_{1}^{*},\cdots,x_{a}^{*})$ is a stable positive equilibrium point of the sub-system of the winners when ignoring the loser agents.
Proof 6.6.
The proof is similar to the proof of Theorem 22.
Remark 6.7.
Since there are multiple sub-systems of (20), each subsystem may have a stable positive equilibrium. In general, multi-stability may occur.
7 Numerical example
@@
Firstly, we conduct simulations for the case of competition between 2 factions. We assume a total of $5$ species, $2$ in one faction and $3$ in the other. Figures 1(a)-1(c) include all the possible results we can observe and are three typical examples for each case. Figure 1(a) is in line with Theorem 18, while Figure 1(b) corresponds to Theorem 20. Similarly, Figure 1(c) reflects the analytical results of Theorem 22. Since all other equilibria are unstable, we don’t observe that the solution converges towards them in the simulations. From figure 1(a) and 1(d), we confirm the bistability properties of the two-faction system. Furthermore, to highlight the influence of HOIs, the system parameters and initial conditions are the same in simulations 1(b) and 1(c) except that we change $d_{1jk}$ to $5d_{1jk}$, respectively. As we increase the $d_{1jk}$, according to the Theorem 22, the first faction dies out, which is indeed in line with the simulation results. This shows that our theory can also be seen as a manipulation strategy to adjust the winner species. Since HOIs usually denote the indirect interaction in ecology, HOIs are potentially more suitable to adjust than pairwise direct interaction. Finally, from simulations, we observe that large self-competition terms ($a_{ii},\hat{a}_{ii},c_{iii},\hat{c}_{iii}$) will improve the chance that the system doesn’t diverge to infinity. This is because diagonally dominant tensors with all positive diagonal entries are $\mathcal{S}$-tensors, which potentially yields a globally stable positive equilibrium.
@@
Then, we use a more general case. For the tensor $A$, its entries are $A_{111}=A_{222}=11,A_{112}=A_{122}=A_{211}=A_{212}=-1,A_{121}=10,A_{221}=5$. For the matrix $B$, $B_{11}=0.8883,B_{12}=1,B_{21}=0,B_{22}=1$. Let $v=(0.1117,1)^{\top}$. We can check that $Av^{2}>\mathbf{0}$ and $Bv>\mathbf{0}$. The tensors $A,B$ are all $\mathcal{S}-$tensors but not $\mathcal{H}^{+}-$ tensors. Thus, the conditions of Theorem 1 are satisfied. Numerical simulation suggests that $Ax^{2}+Bx=\mathbf{1}$ has a unique positive solution $(0.49547,0.43791)^{\top}.$
8 Conclusion
This paper studies a general higher-order Lotka-Volterra model. More precisely, we consider 3 typical scenarios: full cooperation, competition between 2 factions, and pure competition respectively. For the analysis tool, we provide an existence result of a positive equilibrium for a non-homogeneous polynomial system and give a lower and upper bound for the solution of a polynomial complementarity problem under mild conditions, which is an extension of the current results in tensor algebras. Then, we further utilize the properties of tensor and monotone system theory to provide the results regarding the existence, uniqueness, and stability of the equilibrium of the system. This paper yields a comprehensive understanding of a higher-order Lotka-Volterra model. Finally, all theoretical results are illustrated by some numerical examples.